# Approximating Discrimination Within Models When Faced With Several Non-Binary Sensitive Attributes

### v1

@@ 3.1.1 Distance between sets for sensitive attributes with binary values, from our previous work [34]
By recording the true label $y$ and the prediction $\hat{y}$ as one denotation (say $\ddot{y}$) for simplification, we could rewrite Equations (1) and (2) as
$\small\begin{split}\mathbf{D}_{\cdot}(S_{1},\bar{S}_{1})\triangleq\max\big{\{}&\max_{(\bm{x},y)\in S_{1}}\min_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{1}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)},\\ &\max_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{1}}\min_{(\bm{x},y)\in S_{1}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)}\big{\}}\,.\end{split}$ (3)
We will continue using the above notations in the subsequent context for simplification.
3.1.2 Distance between sets for one sensitive attribute with multiple values
As for the scenarios where only one sensitive attribute exists, let $\mathop{\bm{a}}=[a_{i}]^{\mathsf{T}}$ be a single sensitive attribute, in other words, $n_{a}\!=\!1$, $a_{i}\!\in\!\mathcal{A}_{i}\!=\!\{1,2,...,n_{a_{i}}\}$, $n_{a_{i}}\!\geqslant 3$, and $n_{a_{i}}\!\in\!\mathbb{Z}_{+}$. Then the original dataset $S$ can be divided into a few disjoint sets according to the value of this attribute $a_{i}$, that is, $S_{j}=\{(\bm{x},y)\in S\mid a_{i}=j\},\forall j\in\mathcal{A}_{i}$. We can now extend Eq. (3) and introduce the following distance measures: (i) maximal distance measure for one sensitive attribute
@@ 3.1.2 Distance between sets for one sensitive attribute with multiple values
and (ii) average distance measure for one sensitive attribute
$\small\mathbf{D}_{\cdot,\bm{a}}^{\text{avg}}(S,a_{i})\triangleq\frac{1}{n}\sum_{j=1}^{n_{a_{i}}}\sum_{(\bm{x},y)\in S_{j}}\min_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{j}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)}\,,$ (5)
where $\bar{S}_{j}\!=S\!\setminus\!S_{j}$. Notice that $\mathbf{D}_{\cdot,\bm{a}}^{\text{}}(S,a_{i})\!=\!\mathbf{D}_{\cdot}(S_{1},\bar{S}_{1})$ when $\mathcal{A}_{i}\!=\!\{0,1\}$.
3.1.3 Distance between sets for multiple sensitive attributes with multiple values
Now we discuss the general case, where we have several sensitive attributes $\mathop{\bm{a}}=[a_{1},a_{2},...,a_{n_{a}}]^{\mathsf{T}}$ and each $a_{i}\in\mathcal{A}_{i}=\{1,2,..,n_{a_{i}}\}$, where $n_{a_{i}}$ is the number of values for this sensitive attribute $a_{i}\,(1\leqslant i\leqslant n_{a})$. We can now introduce the following generalised distance measures: (i) maximal distance measure for sensitive attributes
@@ 3.1.3 Distance between sets for multiple sensitive attributes with multiple values
and (ii) average distance measure for sensitive attributes
@@
    Both $\mathbf{D}_{\cdot,\bm{a}}^{\text{}}(S,a_{i})$ and $\mathbf{D}_{\cdot,\bm{a}}^{\text{avg}}(S,a_{i})$ measure the fairness regarding the sensitive attribute $a_{i}$.
@@
    As their names suggest, the maximal distance represents the largest possible disparity between instances with different sensitive attributes, while the average distance reflects the average disparity between instances with different sensitive attributes. The formal distance measures are more stringent, they are susceptible to data noise. In contrast, the latter type of distance measures are more resilient against the influence of data noise.
We remark that $\mathbf{D}_{\mathop{\bm{a}}}(S)$, $\mathbf{D}_{\mathop{\bm{a}}}^{\text{avg}}(S)$ reflect the biases from the data and that $\mathbf{D}_{f,\mathop{\bm{a}}}(S)$, $\mathbf{D}_{f,\mathop{\bm{a}}}^{\text{avg}}(S)$ reflect the extra biases from the learning algorithm. Then the following values could be used to reflect the fairness degree of this classifier, that is,
@@
We name the fairness degrees defined as above of one classifier by Eq. (8) as ‘maximum harmonic fairness measure via manifolds (HFM)’ and ‘average HFM’, respectively.
3.2 A prompt approximation of distances between sets for Euclidean spaces
To reduce the high computational complexity ($\mathcal{O}(n^{2})$) of directly calculating Equations (4) and (5), we propose a prompt algorithm that can be viewed as a modification of an $\mathcal{O}(n\log n)$-algorithm in our previous work [34].
We start by recalling the algorithm introduced in [34]. Since the core operation in Equations (4) and (5) is to evaluate the distance between data points inside $\mathcal{X\times Y}$, to reduce the number of distance evaluation operations involved in Eq. (4) and (5), we observe that the distance between similar data points tends to be closer than others after projecting them onto a general one-dimensional linear subspace (refer [34, Lemma 1]). To be concrete, let $g:\mathcal{X\times Y}\mapsto\mathbb{R}$ be a random projection, then we could write $g$ as
@@ 3.2 A prompt approximation of distances between sets for Euclidean spaces
where $\bm{w}\!=\![w_{0},w_{1},...,w_{n_{x}}]^{\mathsf{T}}$ is a non-zero random vector. Now, we choose a random projection $g\!:\!\mathcal{X\!\times\!Y}\!\mapsto\!\mathbb{R}$, then we sort all the projected data points on $\mathbb{R}$. According to [34, Lemma 1], it is likely that for the instance $(\bm{x},y)$ in $S_{j}$, the desired instance $\operatorname*{argmin}_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{j}}\mathbf{d}\big{(}(\breve{\bm{x}},{y}),(\breve{\bm{x}}^{\prime},{y}^{\prime})\big{)}$ would be somewhere near it after the projection, and vice versa. Thus, by using the projections in Eq. (9), we could accelerate the process in Eq. (4) and (5) by checking several adjacent instances rather than traversing the whole dataset.
In this paper, instead of taking one random vector each time, we now take a few orthogonal random vectors each time and do the above process for all these orthogonal vectors. The number of these orthogonal vectors could be $n_{x}+1$, or smaller (such as two or three) if the practitioners would like to save more time in practice. For instance, we set two orthogonal random vectors in Algorithm 2 at present. Then we take the minimum among all estimated distances. This modification may slightly increase the time cost of approximation a bit compared with our previous work [34, Algorithm 1], yet will still significantly accelerate the execution speed and the effectiveness of the projection algorithm, compared with the direct calculation of distances.
Then we could propose an approximation algorithm to estimate the distance between sets in Equations (4) and (5), named as ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’, shown in Algorithm 2. As for the distance in Equations (6) and (7), we propose ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’, shown in Algorithm 1. Note that there exists a sub-route within ApproxDist to obtain an approximated distances between sets, which is named as ‘Acceleration sub-procedure (AcceleDist)’ and shown in Algorithm 3. As the time complexity of sorting in line 2 of Algorithm 3 could reach $\mathcal{O}(n\log n)$, we could get the computational complexity of Algorithm 3 as follows: i) The complexity of line 1 is $\mathcal{O}(n)$; and ii) The complexity from line 4 to line 10 is $\mathcal{O}(2m_{2}+1)$. Thus the overall time complexity of Algorithm 3 would be $\mathcal{O}(n(\log n+m_{2}+1))$, and that of Algorithm 2 be $\mathcal{O}(m_{1}n(\log n+m_{2}))$, and that of Algorithm 1 be $\mathcal{O}(n_{a}m_{1}n(\log n+m_{2}))$. As both $m_{1}$ and $m_{2}$ are the designated constants, and $n_{a}$ is also a fixed constant for one specific dataset, the time complexity of computing the distance is then down to $\mathcal{O}(n\log n)$, which is more welcome than $\mathcal{O}(n^{2})$ for the direct computation in Section 3.1.
@@
0:  Dataset $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}=\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})\}_{i=1}^{n}$ where $\mathop{\bm{a}}_{i}=[a_{i,1},a_{i,2},...,a_{i,n_{a}}]^{\mathsf{T}}$, prediction of $S$ by the classifier $f(\cdot)$ that has been trained, that is, $\{\hat{y}_{i}\}_{i=1}^{n}$, and two hyper-parameters $m_{1}$ and $m_{2}$ as the designated numbers for repetition and comparison respectively
@@
0:  Dataset $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}=\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})\}_{i=1}^{n}$, prediction of $S$ by the classifier $f(\cdot)$ that has been trained, that is, $\{\hat{y}_{i}\}_{i=1}^{n}$, and two hyper-parameters $m_{1}$ and $m_{2}$ as the designated numbers for repetition and comparison respectively
@@
0:  Data points $\{(\mathop{\breve{\bm{x}}}_{i},a_{i})\}_{i=1}^{n}$, its corresponding value $\{\ddot{y}_{i}\}_{i=1}^{n}$, where $\ddot{y}_{i}$ could be its true label $y_{i}$ or prediction $\hat{y}_{i}$ by the classifier $f(\cdot)$, a random vector $\bm{w}$ for projection, and a hyper-parameter $m_{2}$ as the designated number for comparison
@@ 3.3 Algorithmic effectiveness analysis of ApproxDist
We have introduced an important lemma that confirms the observation that ‘the distance between similar data points tends to be closer than others after projecting them onto a general one-dimensional linear subspace’ (refer to [34, Lemma 1]), restated in Lemma 1. It demonstrates by Eq. (10) that the probability $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ also goes to zero when the ratio $\nicefrac{{r_{1}}}{{r_{2}}}$ goes to zero. Additionally, it is easy to observe that $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ reaches the same order of magnitude as $\nicefrac{{r_{1}}}{{r_{2}}}$, and especially, when $r_{1}$ equals $r_{2}$, $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ could be roughly viewed as $\nicefrac{{1}}{{2}}$ for coarse approximation. It means that the breaking probability of the aforementioned statement—similar data points leading to closer distances—tends to increase as $r_{1}$ gradually gets closer to $r_{2}$. And the profound meaning behind Lemma 1 is that the bigger the gap of lengths between $\bm{v}_{1}$ and $\bm{v}_{2}$ is, the more effective and efficient our proposed approximation algorithms would be.
Lemma 1.
@@
Proof.
Notice that $|\langle\bm{w},\bm{v}_{1}\rangle|\geqslant|\langle\bm{w},\bm{v}_{2}\rangle|$ is equivalent to
$\small\langle\bm{v}_{2}-\bm{v}_{1},\bm{w}\rangle\langle\bm{v}_{1}+\bm{v}_{2},\bm{w}\rangle\leqslant 0\,.$ (11)
If $\bm{w}$ satisfies Eq. (11), then it lies between two hyperplanes that are perpendicular to $\bm{v}_{1}-\bm{v}_{2}$ and $\bm{v}_{1}+\bm{v}_{2}$ respectively. Denote by $\theta$ the angle between these two hyperplanes (which is equal to the acute angle between $\bm{v}_{2}-\bm{v}_{1}$ and $\bm{v}_{1}+\bm{v}_{2}$), then $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})=\nicefrac{{\theta}}{{\pi}}$. Moreover,
$\small\sin^{2}\theta=1-\cos^{2}\theta=\frac{4\lVert\bm{v}_{1}\rVert^{2}\lVert\bm{v}_{2}\rVert^{2}-4\langle\bm{v}_{1},\bm{v}_{2}\rangle^{2}}{(\lVert\bm{v}_{1}\rVert^{2}+\lVert\bm{v}_{2}\rVert^{2})^{2}-4\langle\bm{v}_{1},\bm{v}_{2}\rangle^{2}}\,.$ (12)
Here $\lVert\bm{v}_{i}\rVert^{2}=\langle\bm{v}_{i},\bm{v}_{i}\rangle=r_{i}^{2}$ ($i=1,2$) is the square length of the vector. Recall that $\langle\bm{v}_{1},\bm{v}_{2}\rangle=\lVert\bm{v}_{1}\rVert\lVert\bm{v}_{2}\rVert\mathrm{cos}~{}\phi$. By Eq. (12), we have
$\small\frac{\lVert\bm{v}_{1}\rVert^{2}}{\lVert\bm{v}_{2}\rVert^{2}}\sin^{2}\phi\leqslant\sin^{2}\theta\leqslant\frac{4\lVert\bm{v}_{1}\rVert^{2}\lVert\bm{v}_{2}\rVert^{2}}{(\lVert\bm{v}_{1}\rVert^{2}+\lVert\bm{v}_{2}\rVert^{2})^{2}}\,.$ (13)
Combining Eq. (13) with the fact that $\frac{2}{\pi}\theta\leqslant\mathrm{sin}~{}\theta\leqslant\theta$, we conclude that the probability $\mathbb{P}(\bm{v}_{2},\bm{v}_{2})=\frac{\theta}{\pi}$ satisfies the desired inequalities. ∎
Our main result in this subsection is Proposition 2, whereby Eq. (15), the efficiency of ApproxDist decreases as the scaled density $\mu$ of the original dataset increases. Meanwhile, when dealing with large-scale datasets, the more insensitive attributes we have, the more efficient ApproxDist is. In general, the efficiency of ApproxDist depends on the shape of these two subsets of $S$. Roughly speaking, the more separated these two sets are from each other, the more efficient ApproxDist is.
@@
Let $S\!=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}\subset\mathcal{X}\times\mathcal{Y}$ be a $(k\!+\!1)$-dimensional dataset where instances have $(k\!+\!1)$ features, an evenly distributed dataset with a size of $n$ that is a random draw of the feature-label space $\mathcal{X\times Y}$. For any two subsets of $S$ with distance $d$ (ref. Eq. (3)), suppose further that the scaled density
@@
Let $S_{0}$ and $S_{1}$ be two sub-datasets of $S$. We fix the instance $\bm{v}_{0}\in S_{0}$ such that $d\triangleq\mathbf{D}(S_{0},S_{1})=\mathbf{d}(\bm{v}_{0},\bm{v_{1}})$ for some $\bm{v_{1}}\in S_{1}$. For simplicity, we may set $\bm{v}_{0}$ as the origin. The probability that an instance $\bm{v}\in S_{1}$ has a shorter length than $\bm{v}_{1}$ after projection to a line (see Eq. (9)) is denoted as $\mathbb{P}(\bm{v}_{1},\bm{v})$. By assumption, we only need to consider those instances whose length is greater than $\alpha d$ (outside the ball $\mathbf{B}(\alpha d)$ centered at origin). Hence, the desired probability is bounded from below by
$\small 1-\bigg{(}\frac{1}{m_{2}}\sum_{\bm{v}\notin\mathbf{B}(\alpha d)}\mathbb{P}(\bm{v}_{0},\bm{v})\bigg{)}^{m_{1}}\,.$ (16)
@@
Under our assumption, Eq. (17) attains the lowest value when the data are evenly distributed inside a hollow ball $\mathbf{B}_{0}\setminus\mathbf{B}(d)$ centered at $\bm{v}_{0}$. The radius of $\mathbf{B}_{0}$, denoted as $r_{0}$, satisfies
$\small n-1=\mu\frac{\mathrm{Vol}(\mathbf{B}_{0}\setminus\mathbf{B}(d))}{\mathrm{Vol}(\mathbf{B}(d))}=\mu\Big{(}\Big{(}\frac{r_{0}}{d}\Big{)}^{k+1}-1\Big{)}\,.$ (18)

### v2

@@ 3.1.1 Distance between sets for one bi-valued SA, from our previous work [36]
For simplification, we could rewrite Equations (1) and (2) as
$\small\begin{split}\mathbf{D}_{\cdot}(S_{1},\bar{S}_{1})\triangleq\max\big{\{}&\max_{(\bm{x},y)\in S_{1}}\min_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{1}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)},\\ &\max_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{1}}\min_{(\bm{x},y)\in S_{1}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)}\big{\}}\,,\end{split}$ (3)
by recording the true label $y$ and the prediction $\hat{y}$ as one denotation (say $\ddot{y}$). We will continue using the above notations in the subsequent context.
3.1.2 Distance between sets for one multi-valued SA
As for the scenarios where only one SA exists, let $\mathop{\bm{a}}=[a_{i}]^{\mathsf{T}}$ be a single SA, in other words, $n_{a}\!=\!1$, $a_{i}\!\in\!\mathcal{A}_{i}\!=\!\{1,2,...,n_{a_{i}}\}$, $n_{a_{i}}\!\geqslant 3$, and $n_{a_{i}}\!\in\!\mathbb{Z}_{+}$. Then the original dataset $S$ can be divided into a few disjoint sets according to the value of this attribute $a_{i}$, that is, $S_{j}=\{(\bm{x},y)\in S\mid a_{i}=j\},\forall j\in\mathcal{A}_{i}$. We can now extend Eq. (3) and introduce the following distance measures: (i) maximal distance measure for one SA,
@@ 3.1.2 Distance between sets for one multi-valued SA
and (ii) average distance measure for one SA,
$\small\textstyle\mathbf{D}_{\cdot,\bm{a}}^{\text{avg}}(S,a_{i})\!\triangleq\!\frac{1}{n}\sum_{j=1}^{n_{a_{i}}}\sum_{(\bm{x},y)\in S_{j}}\min_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{j}}\mathbf{d}\big{(}(\breve{\bm{x}},{\ddot{y}}),(\breve{\bm{x}}^{\prime},{\ddot{y}}^{\prime})\big{)}\,,$ (5)
where $\bar{S}_{j}\!=S\!\setminus\!S_{j}$. Notice a special case that $\mathbf{D}_{\cdot,\bm{a}}^{\text{}}(S,a_{i})\!=\!\mathbf{D}_{\cdot}(S_{1},\bar{S}_{1})$ when $\mathcal{A}_{i}\!=\!\{0,1\}$.
3.1.3 Distance between sets for several multi-valued SAs
Now we discuss the general case, where we have several SAs $\mathop{\bm{a}}\!=\![a_{1},a_{2},...,a_{n_{a}}]^{\mathsf{T}}$ and each $a_{i}\!\in\!\mathcal{A}_{i}\!=\!\{1,2,..,n_{a_{i}}\}$, where $n_{a_{i}}$ is the number of values for this SA $a_{i}\,(1\!\leqslant\!i\!\leqslant\!n_{a})$. We can now introduce the following generalised distance measures: (i) maximal distance measure for multiple SAs,
@@ 3.1.3 Distance between sets for several multi-valued SAs
and (ii) average distance measure for multiple SAs,
@@
    Both $\mathbf{D}_{\cdot,\bm{a}}^{\text{}}(S,a_{i})$ and $\mathbf{D}_{\cdot,\bm{a}}^{\text{avg}}(S,a_{i})$ measure the fairness regarding the SA $a_{i}$.
@@
    As their names suggest, the maximal distance represents the largest possible disparity between instances with different SAs, while the average distance reflects the average disparity between instances with different SAs. The formal distance measures are more stringent, they are susceptible to data noise. In contrast, the latter type of distance measures are more resilient against the influence of data noise.
We remark that $\mathbf{D}_{\mathop{\bm{a}}}(S)$, $\mathbf{D}_{\mathop{\bm{a}}}^{\text{avg}}(S)$ reflect the biases from the data and that $\mathbf{D}_{f,\mathop{\bm{a}}}(S)$, $\mathbf{D}_{f,\mathop{\bm{a}}}^{\text{avg}}(S)$ reflect the biases from the learning algorithm. Then, the following values could be used to capture the intensity of variation between them, i.e.
@@
That is to say, we use them to capture the extra discrimination introduced in the learning procedure, reflecting the fairness degree of this classifier. We name the fairness degrees defined as above of one classifier by Eq. (8) as ‘maximum Harmonic Fairness measure via Manifolds (HFM)’ and ‘average HFM’, respectively.
3.2 An efficient approximation of distances between sets for Euclidean spaces
To reduce the high computational complexity ($\mathcal{O}(n^{2})$) of directly calculating Equations (4) and (5),444 The original time complexity of the direct computation in Eq. (3) given a dataset of size $n$ is $\mathcal{O}(2n_{1}(n-n_{1}))$ where $n_{1}$ is the size of $S_{1}$, that is, $\mathcal{O}(n^{2})$. By extension, the time complexity of the direct computation in Eq. (4) and (5) is $\mathcal{O}(\sum_{j=1}^{n_{a_{i}}}n_{j}(n-n_{j}))$ where $n_{j}$ is the size of $S_{j}$ for all $j\in\{1,2,...,n_{a_{i}}\}$, that is, $\mathcal{O}(n^{2})$ as well. we propose an efficient algorithm that can be viewed as a modification of an $\mathcal{O}(n\log n)$-algorithm in our previous work [36].
We start by recalling the algorithm introduced in [36]. Since the core operation in Equations (4) and (5) is to evaluate the distance between data points inside $\mathcal{X\times Y}$, to reduce the number of distance evaluation operations involved in Eq. (4) and (5), we observe that the distance between similar data points tends to be closer than others after projecting them onto a general one-dimensional linear subspace (refer [36, Lemma 1]). To be concrete, let $g:\mathcal{X\times Y}\mapsto\mathbb{R}$ be a random projection, then we could write $g$ as
@@ 3.2 An efficient approximation of distances between sets for Euclidean spaces
where $\bm{w}\!=\![w_{0},w_{1},...,w_{n_{x}}]^{\mathsf{T}}$ is a non-zero random vector. Now, we choose a random projection $g\!:\!\mathcal{X\!\times\!Y}\!\mapsto\!\mathbb{R}$, then we sort all the projected data points on $\mathbb{R}$. According to [36, Lemma 1], it is likely that for the instance $(\bm{x},y)$ in $S_{j}$, the desired instance $\operatorname*{argmin}_{(\bm{x}^{\prime},y^{\prime})\in\bar{S}_{j}}\mathbf{d}\big{(}(\breve{\bm{x}},{y}),(\breve{\bm{x}}^{\prime},{y}^{\prime})\big{)}$ would be somewhere near it after the projection, and vice versa. Thus, by using the projections in Eq. (9), we could accelerate the process in Eq. (4) and (5) by checking several adjacent instances rather than traversing the whole dataset.
In this paper, instead of taking one random vector each time, we now take a few orthogonal random vectors each time and do the above process for all these orthogonal vectors. The number of these orthogonal vectors could be $n_{x}+1$, or smaller (such as two or three), if the practitioners would like to save more time in practice. For instance, we set two orthogonal random vectors [^5] In practice, we first generate a ($1+n_{x}$)-dimensional matrix randomly (for example, from the standard normal distribution), and then use the QR decomposition to get $Q$—an orthogonal matrix. The columns of $Q$ are orthogonal unit vectors, and the first two (or three, depending on the number of orthogonal vectors that the practitioner likes) columns will be used as the orthogonal vectors in line 2 of Algorithm 2. in Algorithm 2 at present. Then we take the minimum among all estimated distances. This modification may slightly increase the time cost of approximation a bit compared with our previous work [36, Algorithm 1], yet will still significantly accelerate the execution speed and the effectiveness of the projection algorithm, compared with the direct calculation of distances.
@@
0:  Dataset $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}=\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})\}_{i=1}^{n}$ where $\mathop{\bm{a}}_{i}=[a_{i,1},a_{i,2},...,a_{i,n_{a}}]^{\mathsf{T}}$, prediction of $S$ by the classifier $f(\cdot)$ that has been trained, that is, $\{\hat{y}_{i}\}_{i=1}^{n}$, and two hyperparameters $m_{1}$ and $m_{2}$ as the designated numbers for repetition and comparison respectively
@@
0:  Dataset $S=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}=\{(\mathop{\breve{\bm{x}}}_{i},\mathop{\bm{a}}_{i},y_{i})\}_{i=1}^{n}$, prediction of $S$ by the classifier $f(\cdot)$ that has been trained, that is, $\{\hat{y}_{i}\}_{i=1}^{n}$, and hyperparameters $m_{1}$, $m_{2}$ as the designated number for repetition and comparision
@@
0:  Data points $\{(\mathop{\breve{\bm{x}}}_{i},a_{i})\}_{i=1}^{n}$, its corresponding value $\{\ddot{y}_{i}\}_{i=1}^{n}$, where $\ddot{y}_{i}$ could be its true label $y_{i}$ or prediction $\hat{y}_{i}$ by the classifier $f(\cdot)$, a random vector $\bm{w}$ for projection, and a hyperparameter $m_{2}$ as the designated number for comparison
@@
Then we could propose an approximation algorithm to estimate the distance between sets in Equations (4) and (5), named as ‘Approximation of distance between sets for one sensitive attribute with multiple values (ApproxDist)’, shown in Algorithm 2. As for the distance in Equations (6) and (7), we propose ‘Approximation of extended distance between sets for several sensitive attributes with multiple values (ExtendDist)’, shown in Algorithm 1. Note that there exists a sub-route within ApproxDist to obtain an approximated distance between sets, which is named as ‘Acceleration sub-procedure (AcceleDist)’ and shown in Algorithm 3. As the time complexity of sorting in line 2 of Algorithm 3 could reach $\mathcal{O}(n\log n)$, we could get the computational complexity of Algorithm 3 as follows: i) The complexity of line 1 is $\mathcal{O}(n)$; and ii) The complexity from line 4 to line 10 is $\mathcal{O}(2m_{2}+1)$. Thus the overall time complexity of Algorithm 3 would be $\mathcal{O}(n(\log n+m_{2}+1))$, and that of Algorithm 2 be $\mathcal{O}(m_{1}n(\log n+m_{2}))$, and that of Algorithm 1 be $\mathcal{O}(n_{a}m_{1}n(\log n+m_{2}))$. As both $m_{1}$ and $m_{2}$ are the designated constants, and $n_{a}$ is also a fixed constant for one specific dataset, the time complexity of computing the distance is then down to $\mathcal{O}(n\log n)$, which is more welcome than $\mathcal{O}(n^{2})$ for the direct computation in Section 3.1.
@@ 3.3 Algorithmic effectiveness analysis of ApproxDist
In higher dimensional statistics, it is known that if we take a random projection $A$ from a n-dimensional space to a $k$-dimensional subspace, then for every vector $\mathbf{x}$, we have the following probability estimation:
$\small\mathbb{P}\left(\left|\frac{|A\mathbf{x}|}{|\mathbf{x}|}-\sqrt{\frac{k}{n}}\right|\leq\varepsilon\sqrt{\frac{k}{n}}\right)\geq 1-C\exp\{-Ck\varepsilon^{2}\}\,,$ (10)
where $C$ is a universal constant independent of the dimensional $n,k$ and the projection $A$. In particular, when $k=1$, we have proved a simplified version of (10) (refer to [36, Lemma 1]), restated in Lemma 1. These results support the following observation: ‘the distance between similar data points tends to be closer than others after projecting them onto a general one-dimensional linear subspace.’ It demonstrates by Eq. (11) that the probability $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ also goes to zero when the ratio $\nicefrac{{r_{1}}}{{r_{2}}}$ goes to zero. Additionally, it is easy to observe that $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ reaches the same order of magnitude as $\nicefrac{{r_{1}}}{{r_{2}}}$, and especially, when $r_{1}$ equals $r_{2}$, $\mathbb{P}(\bm{v}_{1},\bm{v}_{2})$ could be roughly viewed as $\nicefrac{{1}}{{2}}$ for coarse approximation. It means that the breaking probability of the aforementioned statement—similar data points leading to closer distances—tends to increase as $r_{1}$ gradually gets closer to $r_{2}$. And the profound meaning behind Lemma 1 is that the bigger the gap of lengths between $\bm{v}_{1}$ and $\bm{v}_{2}$ is, the more effective and efficient our proposed approximation algorithms would be.
Lemma 1 (Lemma 1 [36]).
@@
Our main result in this subsection is Proposition 2 (a modified version of [36, Proposition 2]), whereby Eq. (13), the efficiency of ApproxDist decreases as the scaled density $\mu$ of the original dataset increases. Meanwhile, when dealing with large-scale datasets, the more insensitive attributes we have, the more efficient ApproxDist is. In general, the efficiency of ApproxDist depends on the shape of these two subsets of $S$. Roughly speaking, the more separated these two sets are from each other, the more efficient ApproxDist is.
@@
Let $S\!\!=\!\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}\!\subset\!\mathcal{X}\!\times\!\mathcal{Y}$ be a $(k\!+\!1)$-dimensional dataset where instances have $(k\!+\!1)$ features, an evenly distributed dataset with a size of $n$ that is a random draw of the feature-label space $\mathcal{X\!\times\!Y}$. For any two subsets of $S$ with distance $d$ (ref. Eq. (3) and (4)), suppose further that the scaled density
@@
Let $S_{1},S_{2},...,S_{n_{a_{i}}}$ be $n_{a_{i}}$ sub-datasets of $S$. For each $j\in\{1,2,...,n_{a_{i}}\}$, then $S_{j}$ and $\bar{S}_{j}$ be two sub-datasets of $S$. We fix the instance $\bm{v}_{j}\in S_{j}$ such that $d\triangleq\mathbf{D}_{\mathop{\bm{a}}}(S,a_{i})=\mathbf{d}(\bm{v}_{j},\bm{v}_{k})$ for some $\bm{v}_{k}\in\bar{S}_{j}$. For simplicity, we may set $\bm{v}_{j}$ as the origin. The probability that an instance $\bm{v}\in\bar{S}_{j}$ has a shorter length than $\bm{v}_{k}$ after projection to a line (see Eq. (9)) is denoted as $\mathbb{P}(\bm{v}_{k},\bm{v})$. By assumption, we only need to consider those instances whose length is greater than $\alpha d$ (outside the ball $\mathbf{B}(\alpha d)$ centred at the origin). Hence, the desired probability is bounded from below by
$\small 1-\bigg{(}\frac{1}{m_{2}}\sum_{\bm{v}\notin\mathbf{B}(\alpha d)}\mathbb{P}(\bm{v}_{j},\bm{v})\bigg{)}^{m_{1}}\,.$ (14)
@@
Under our assumption, Eq. (15) attains the lowest value when the data are evenly distributed inside a hollow ball $\mathbf{B}_{j}\setminus\mathbf{B}(d)$ centred at $\bm{v}_{j}$. The radius of $\mathbf{B}_{j}$, denoted as $r_{j}$, satisfies
$\small n-1=\mu\frac{\mathrm{Vol}(\mathbf{B}_{j}\setminus\mathbf{B}(d))}{\mathrm{Vol}(\mathbf{B}(d))}=\mu\Big{(}\Big{(}\frac{r_{j}}{d}\Big{)}^{k+1}-1\Big{)}\,.$ (16)